# Eco-evolutionary constraints for the endemicity of rapidly evolving viruses

### v1

@@ Discussion
Our findings should be considered in light of several limitations of the minimal framework here introduced. Biologically, the gradual evolution of the virus is challenged by the reported anomalous diffusion behaviors of viruses across the genomic space [57] or the complex genotype-phenotype networks shaping the evolution of viruses [58]. The evolutionary dynamics in our model could also be refined by introducing trade-off mechanisms [59], coupling the evolution of different epidemiological parameters. Despite the scarce empirical evidence supporting them [60], evolutionary trade-offs are typically included in models to warrant the existence of evolutionary stable strategies [61]. Our first exploration with the recovery-transmissibility trade-off shows an uneven impact of such mechanism, keeping our findings for some viruses but altering them for others. At the population level, assuming all the individuals to be equivalent does not allow our model to capture the strong influence reported for superspreaders [62, 63] or immuno-compromised patients [33, 64] on the eco-evolutionary dynamics of viruses. Notwithstanding all these limitations, the generality of our findings, based on the relationship between epidemiological and evolutionary time scales, make us feel confident about their ubiquity in more biologically-grounded models.
@@
In a nutshell, our results underscore the relevance of the interplay between different evolutionary pathways of rapidly evolving virus in shaping their associated epidemic trajectories. Improving our knowledge on this topic will enhance our preparedness against future epidemic threats and the use of eco-evolutionary frameworks as reliable benchmarks to predict the evolution of emergent pathogens and design optimal control policies to mitigate both their short-term and long-term impact on society.
Materials and Methods
Stochastic simulations
@@
At the first stage, we assume that infected individuals recover with a probability $Pr(I\rightarrow R)=1-e^{-\mu\Delta t}$. To simulate contagion processes, we assume that each susceptible agent $i$ chooses $k$ neighbors at random and compute the probability of infection as $Pr(S_{i}\rightarrow I)=1-e^{-\sum\limits_{l\in\Gamma(i)}\lambda_{l}\Delta t}$, where $\Gamma(i)$ denotes the set of infected neighbors of agent $i$ at that time step. If a contagion event is successful, the variant contracted by the agent is chosen proportionally to the infection rates of the variants existing in $\Gamma(i)$. Therefore, the probability that agent $i$ contracts the disease from an infected neighboring agent $j$ is $Pr(S_{i}\rightarrow I_{j})=\dfrac{\lambda_{j}}{\sum\limits_{l\in\Gamma(i)}\lambda_{l}}Pr(S_{i}\rightarrow I)$. For reinfection processes, the same rationale to choose the variant applies but now considering that the individual rates at which a recovered agent $i$ contracts the disease from an infected agent $j$ is $\lambda^{\prime}_{ij}=\Theta(x_{j}-x_{i})\lambda_{j}\left(1-e^{-(x_{j}-x_{i})}\right)$, where $\Theta(x)$ is the Heavyside function introduced to avoid reinfection from past variants.
@@
Once the epidemiological state of each individual is updated, evolution of the virus only occurs inside infected individuals. Therefore, the transmissibility (antigenic position) of an agent $j$ variant $\lambda_{j}$ ($x_{j}$) changes according to $\lambda_{j}(t+\Delta t)=\lambda_{j}+D_{\lambda}\Delta t\mathcal{N}(0,1)$ $\left(x_{j}(t+\Delta t)=x_{j}+D_{x}\Delta t\mathcal{N}(0,1)\right)$, where $\mathcal{N}(0,1)$ represents a random number drawn from the standard normal distribution.
Approximation to the case reproduction number $\mathcal{R}_{case}(t)$
@@
In the context of an epidemic outbreak, $\mathcal{R}_{case}(t)$ computes the expected number of contagions made by individuals becoming infected at time $t$. The exact computation of such quantity requires knowing the shape of the distribution of variants in both the transmissibility and antigenic spaces. For the sake of simplicity, we opt for providing a rough estimation of the case reproduction number, $\mathcal{R}^{app}_{case}(t)$. For such purpose, we assume that the standard deviation of the aforementioned distribution $\sigma_{x}$ represents the typical distance observed in the antigenic space and that the average transmissibility $\langle\lambda\rangle$ also constitutes a representative quantity for contagion dynamics. Moreover, let us also neglect the changes in the pool of susceptible and recovered population and the evolution of the virus during the infectious period of the focal agent. Under these assumptions, the approximated case reproduction number can be readily computed as:
$\mathcal{R}^{app}_{case}(t)=\frac{\langle\lambda\rangle k}{\mu}\left[\rho_{S}(t)+\rho_{R}(t)\left(1-e^{-\sigma_{x}}\right)\right]\ ,$ (5)
@@
Epidemic prevalence at equilibrium for a virus with deterministic antigenic escape
In this section, we derive the epidemic prevalence at the endemic equilibrium for a virus evolving deterministically its antigenic position and not evolving its transmissibility. For such purpose, let us first define $\rho_{I}(x,t)$ ($\rho_{R}(x,t)$) as the fraction of the population in the infected (recovered) state carrying the variant $x$ at time $t$. Considering antigenic escape and reinfection events, the time evolution of the latter is given by [19, 32]:
$\frac{d\rho_{R}(x,t)}{dt}$ $=\mu\rho_{I}(x,t)$
$-\lambda k\rho_{R}(x,t)\int\rho_{I}(x^{\prime},t)\left(1-e^{-(x^{\prime}-x)}dx^{\prime}\right)$ (6)
Assuming a deterministic evolution of the virus greatly simplifies the set of equations governing the unfolding of the epidemic trajectories, as there is a single circulating strain in the population. Mathematically, this implies that $x_{c}(t)=\widetilde{D}_{x}t$ for the circulating pathogen and hence $\rho_{I}(x,t)=\rho_{I}(t)\delta\left(x-x_{c}(t)\right)$, where $\delta$ represents the Dirac delta function. Rewriting Eq. 7 using a travelling wave approach, we obtain:
@@
where $\tau$ refers to the time elapsed since the recovery of the agent. Assuming that the disease has reached the endemic equilibrium, i.e. $\rho_{I}(0)=\rho_{I}^{*}$, and a regime of strong immune pressure, i.e. $\widetilde{D}_{x}\tau\ll 1$, we solve the previous ODE obtaining:
$\rho_{R}(\tau)=\mu\rho_{I}^{*}e^{-\widetilde{D}_{x}\lambda k\rho_{I}^{*}\tau^{2}/2}\ ,$ (8)
[…]
$\sqrt{\frac{\mu\pi\rho_{I}^{*}}{2\mathcal{R}_{0}\widetilde{D}_{x}}}+\rho_{I}^{*}=1\ .$ (9)
@@
Assuming a very small prevalence in the endemic regime, i.e. $\rho_{I}^{*}\ll 1$, we can estimate $\rho_{I}^{*}$ as
$\rho_{I}^{*}\simeq\frac{2\mathcal{R}_{0}\widetilde{D}_{x}}{\mu\pi}\ .$ (10)
Time for an epidemic peak in a SIR model
@@
For a well-mixed population with constant size, the exact time to the epidemic peak $t_{peak}$ in the standard SIR model can be approximated as [39]:
@@
where $i_{0}$ refers to the fraction of population initially infected. This expression represents a first-order approximation to the exact computation of the epidemic peak in the standard SIR model. Nonetheless, it provides reasonably accurate estimations of the peak in most of the epidemic scenarios [39]. In our model, the latter expression is used to obtain a characteristic time scale during which viruses can evolve their transmissibility to avoid getting extinct. Consequently, note that the former expression comes with intrinsic limitations, as it assumes a fixed value of the infectiousness $\lambda$ and not a evolving quantity over time. Therefore, we should expect predictions to be misaligned with the actual epidemic peak for quickly evolving viruses with low initial infectiousness, i.e. ${\mathcal{R}_{0}\simeq 1}$. In that region, the slow propagation of the virus leads to a substantial evolution in the virus infectiousness throughout the first epidemic wave, which in turn produces a much shorter epidemic wave than the one predicted by Eq. 11.

### v2

@@ Discussion
Our findings should be considered in light of several limitations of the minimal framework here introduced. To save computation time, we have implemented discrete-time simulations with synchronous updates rather than using Gillespie-like algorithms. Biologically, the gradual evolution of the virus is challenged by the reported anomalous diffusion behaviors of viruses across the genomic space [58] or the complex genotype-phenotype networks shaping the evolution of viruses [59]. The evolutionary dynamics in our model could also be refined, as we consider that evolution in infectiousness and antigenic position is unbounded and our analytical results rely on a constant evolution in viral traits. To bound the values of the epidemiological parameters, one can explore the effect of different trade-off mechanisms [60], coupling their joint evolution. Despite the scarce empirical evidence supporting them [61], evolutionary trade-offs are typically included in models to warrant the existence of evolutionarily stable strategies [62]. Our first exploration with the recovery-transmissibility trade-off shows an uneven impact of such mechanism, keeping our findings for some viruses but altering them for others. At the population level, assuming all individuals to be equivalent does not allow our model to capture the strong influence reported for superspreaders [63, 64] or immuno-compromised patients [33, 65] on the eco-evolutionary dynamics of viruses. Moreover, the well-mixed assumption can be improved to include spatially distributed population, as host mobility has been reported to strongly shape the antigenic escape mechanism [66]. Notwithstanding all these limitations, the generality of our findings, based on the relationship between epidemiological and evolutionary time scales, makes us feel confident about their ubiquity in more biologically-grounded models.
@@
In a nutshell, our results underscore the relevance of the interplay between different evolutionary pathways of rapidly evolving viruses in shaping their associated epidemic trajectories. Improving our knowledge on this topic will enhance our preparedness against future epidemic threats and the use of eco-evolutionary frameworks as reliable benchmarks to predict the evolution of emergent pathogens and design optimal control policies to mitigate both their short-term and long-term impact on society.
APPENDIX A: Stochastic simulations
@@
At the first stage, we assume that infected individuals recover with a probability $Pr(I\rightarrow R)=1-e^{-\mu\Delta t}$. To simulate contagion processes, we assume that each susceptible agent $i$ chooses $k$ neighbors at random and compute the probability of infection as $Pr(S_{i}\rightarrow I)=1-e^{-\sum\limits_{l\in\Gamma(i)}\lambda_{l}\Delta t}$, where $\Gamma(i)$ denotes the set of infected neighbors of agent $i$ at that time step. If a contagion event is successful, the variant contracted by the agent is chosen proportionally to the infectiousness of the variants existing in $\Gamma(i)$. Therefore, the probability that agent $i$ contracts the disease from an infected neighboring agent $j$ is $Pr(S_{i}\rightarrow I_{j})=\dfrac{\lambda_{j}}{\sum\limits_{l\in\Gamma(i)}\lambda_{l}}Pr(S_{i}\rightarrow I)$. For reinfection processes, the same rationale to choose the variant applies, but now considering that the individual rates at which a recovered agent $i$ contracts the disease from an infected agent $j$ is $\lambda^{\prime}_{ij}=\Theta(x_{j}-x_{i})\lambda_{j}\left(1-e^{-(x_{j}-x_{i})}\right)$, where $\Theta(x)$ is the Heaviside function introduced to avoid reinfection from past variants.
@@
Once the epidemiological state of each individual is updated, evolution of the virus only occurs inside infected individuals. Therefore, the infectiousness (antigenic position) of an infected agent $j$ $\lambda_{j}$ ($x_{j}$) changes according to $\lambda_{j}(t+\Delta t)=\lambda_{j}+D_{\lambda}\Delta t\mathcal{N}(0,1)$ $\left(x_{j}(t+\Delta t)=x_{j}+D_{x}\Delta t\mathcal{N}(0,1)\right)$, where $\mathcal{N}(0,1)$ represents a random number drawn from the standard normal distribution. Conversely, the antigenic position of recovered individuals remains constant.
APPENDIX B: Interplay between evolution in infectiousness, immune escape and viral endemicity.
In this section, we extend the results reported in Figure 3 about the influence of the stochastic evolution of both antigenic and non-antigenic traits on epidemic outbreaks. We start by considering a fixed value of the speed of evolution in antigenic escape ($D_{x}=0.015$) and different speeds of evolution in infectiousness $D_{\lambda}$. Figs 6A-B confirm that evolution in infectiousness is accelerated at early stages of the outbreak for pathogens with low initial infectiousness. As stated in the main manuscript, this acceleration is more pronounced for high speeds of evolution (Fig 6B), giving rise to an earlier collapse of the curves associated to different infectiousness of the wild-type variants.
We are also interested in determining how evolution in infectiousness shapes the dependence of virus endemicity on the reproduction number of the wild-type variant. Figs. 6C-D show that, for $D_{x}=0.015$, the non-monotonic behavior penalizing the endemicity of viruses with intermediate infectiousness appears consistently regardless of the speed of evolution in infectiousness. When antigenic escape gains relevance, i.e. for large $D_{x}$ values, we instead observe how the non-monotonic behavior disappears as viruses evolve more quickly, yielding the usual increase in endemicity with the infectiousness of the wild-type variant. These results show that changing their intrinsic infectiousness represents an alternative evolutionary pathway for viruses to partially overcome existing immune pressure in the population.
Figure 6: Eco-evolutionary dynamics under stochastic evolution of antigenic and non-antigenic traits. A)-B): Time evolution of the basic reproduction number ${\cal R}_{0}$ for endemic epidemic outbreaks. Line color denotes the basic reproduction number of the wild-type variant ${\cal R}_{0}^{wt}$. The symbol $\langle\cdot\rangle$ denotes that each curve is the result of averaging the individual curves of all endemic realizations observed after simulating 10001000 epidemic outbreaks for each ${\cal R}^{wt}_{0}$ value. The speed of evolution in the antigenic space is set to $D_{x}=0.015$. C-D): Endemicity $f_{endemic}$ of the virus as a function of the basic reproduction number of the wild-type variant ${\cal R}_{0}^{wt}$. Line color denotes the speed of evolution in the antigenic space $D_{x}$. In the panels, the values of the speed of evolution in infectiousness used to simulated epidemics are: (A,C) $D_{\lambda}=0.0004$ and (B,D) $D_{\lambda}=0.0006$. In all the panels, the rest of epidemiological parameters are the same as in Fig. 2 of the main manuscript.
APPENDIX C: Approximation to the case reproduction number $\mathcal{R}_{case}(t)$
@@
In the context of an epidemic outbreak, $\mathcal{R}_{case}(t)$ computes the expected number of contagions made by individuals becoming infected at time $t$. The exact computation of this quantity would require knowing the shape of the distribution of variants in both the infectiousness and antigenic spaces. For the sake of simplicity, we opt for providing a rough estimate of the case reproduction number, $\mathcal{R}^{app}_{case}(t)$. For this purpose, we assume that the standard deviation $\sigma_{x}$ represents a typical distance observed in the antigenic space and that the mean infectiousness $\bar{\lambda}$ of the circulating variants also constitutes a representative quantity for contagion dynamics. Moreover, let us also neglect the changes in the pool of susceptible and recovered population and the evolution of the virus during the infectious period of the focal agent. Under these assumptions, the approximated case reproduction number can be readily computed as:
$\mathcal{R}^{app}_{case}(t)=\frac{\bar{\lambda}}{\mu}\left[\rho_{S}(t)+\rho_{R}(t)\left(1-e^{-\sigma_{x}}\right)\right]\ ,$ (7)
@@
APPENDIX D: Transmission-recovery tradeoff
The model presented in the main manuscript does not include any biological trade-off linking the different epidemiological parameters of pathogens and, therefore, their evolution. These trade-offs are widely adopted in the modelling community to produce evolutionarily stable strategies (ESS) bounding the evolution of viruses. Among them, the virulence-transmissibility trade-off is undoubtely the most studied one following the seminal work by Anderson & May [67]. The virulence-transmissibility trade-off assumes that the virulence of a virus, defined as the rate at which the host die because of the pathogen, increases with the infectiousness of the virus. This increased lethality shortens the infectious period during which the virus host can transmit the pathogen to other individuals. This double-edge sword creates an optimal strategy at intermediate infectiousness, maximizing the expected size of the offspring produced by infected individuals in next generations.
As our framework does not incorporate any birth-death processes involving hosts, we opt for not including the virulence-transmissibility trade-off. Instead, we can study how the joint evolution in infectiousness and in the infectious period affects the eco-evolutionary dynamics of pathogens by introducing the so-called transmission-recovery trade-off [38]. This trade-off implies that the infectious period is shortened for highly transmissible pathogens and emerges naturally when both the production of immune defenses against a pathogen and its inter-host infectiousness are assumed to be proportional to the pathogen load inside hosts.
To account for the transmission-recovery trade-off, we decompose the recovery rate of individuals $\mu$ into two components: the baseline recovery rate $\mu_{0}$, constant and independent of the pathogen load, and the recovery rate $\mu^{\prime}$ changing over time as a result of the evolution in infectiousness $\lambda$ and the aforementioned trade-off. Mathematically, we assume $\lambda=\gamma\sqrt{\mu^{\prime}}$, where $\gamma$ is a constant factor related to the basic reproduction number of the wild-type variant ${\cal R}^{wt}_{0}$ and the initial value of the evolving recovery rate $\mu^{\prime}_{(t=0)}$. Namely:
$\gamma=\frac{{\cal R}_{0}^{wt}\left(\mu_{0}+\mu^{\prime}_{(t=0)}\right)}{k\sqrt{\mu^{\prime}_{(t=0)}}}\ .$ (8)
Let us first consider viruses with ${\cal R}_{0}^{wt}=2$ evolving both their infectiousness ($D_{\lambda}=0.0003$) and their antigenic position ($D_{x}=0.015$). To analyze the impact of the transmission-recovery trade-off, we fix the baseline recovery rate to $\mu_{0}=1.0/7$ days-1 and study the impact of varying the initial value of the evolving recovery rate, $\mu^{\prime}_{(t=0)}$, on the epidemic curves observed in the population. When the transmission-recovery trade-off plays a minor role, i.e. when $\mu^{\prime}_{(t=0)}\ll\mu_{0}$, the epidemic curves represented in Fig. 7A resemble those shown in Fig. 3, characterized by a first prominent epidemic wave followed by an epidemic bottleneck and an endemic regime with a steady growth rate in the number of cases. However, as both rates become comparable, the trade-off bounds the evolution of the virus and considerably slows down the growth rate of cases in the endemic regime, yielding a roughly constant prevalence of the disease for some epidemic scenarios. This deceleration also affects the evolution of the basic reproduction number $\mathcal{R}_{0}$}, which is slower as the trade-off gains more relevance (Fig. 6B).
Figure 7: Impact of the transmission-recovery trade-off on eco-evolutionary dynamics. A)-B): Time evolution of different epidemiological quantities and virus traits. The quantities shown correspond to: (A) fraction of infected population $\rho_{I}$ and (B) basic reproduction number ${\cal R}_{0}$. The symbol $\langle\cdot\rangle$ denotes that each curve is the result of averaging the individual curves of all endemic realizations observed after simulating 10001000 epidemic outbreaks for each ${\cal R}_{0}$ value. The speeds of evolution in the infectiousness and antigenic spaces are set to $D_{\lambda}=0.0003$ and $D_{x}=0.015$ respectively. C): Endemicity $f_{endemic}$ of the virus as a function of the basic reproduction number of the wild-type variant ${\cal R}_{0}^{wt}$. In all the panels, line color denotes the initial value of the evolving recovery rate $\mu^{\prime}_{t=0}$ and the rest of epidemiological parameters are the same as in Fig. 2.
The transmission-recovery trade-off alters the epidemiological time scales of evolving viruses by shortening the infectious period of infected individuals. This phenomenon also changes the features of the viruses more prone to persist in populations (Fig. 7C). Namely, the non-mononotic behavior of viral endemicity with ${\cal R}_{0}^{wt}$ is preserved when $\mu^{\prime}_{(t=0)}\ll\mu_{0}$. Conversely, when $\mu^{\prime}_{(t=0)}\simeq\mu_{0}$, viral endemicity always increases with the basic reproduction number of the wild-type variant. We believe this phenomenon arises from the shortening of the infectious period and the slowing down in the evolution of infectiousness, making it difficult for weakly infectious pathogens to reach the critical values of the basic reproduction number needed to become endemic.
@@
Appendix E: Epidemic prevalence at equilibrium with constant antigenic escape
@@
As shown in the main manuscript, assuming a constant evolution of the antigenic position allows getting an ODE, Eq. (3) governing the time evolution of the density of recovered population $\rho_{R}(\tau)$, where $\tau$ refers to the time elapsed since the recovery of an agent. Considering a steady prevalence of the disease at the epidemic bottleneck, i.e. $\rho_{I}(0)=\rho_{I}^{bottleneck}$, and a regime of strong immune pressure, i.e. $\widetilde{D}_{x}\tau\ll 1$, we solve the previous ODE obtaining:
$\rho_{R}(\tau)=\mu\rho_{I}^{bottleneck}e^{-\widetilde{D}_{x}\lambda k\rho_{I}^{bottleneck}\tau^{2}/2}\ ,$ (9)
where we have used $\lim_{\tau\to 0}\rho_{R}(\tau)=\mu\rho_{I}^{bottleneck}$. To find the epidemic prevalence, we should recall that no susceptible individuals remain in the endemic equilibrium and use the normalization condition $\int\rho_{R}(\tau)d\tau+\rho_{I}^{bottleneck}=1$, yielding:
$\sqrt{\frac{\mu\pi\rho_{I}^{bottleneck}}{2\mathcal{R}_{0}\widetilde{D}_{x}}}+\rho_{I}^{bottleneck}=1\ .$ (10)
@@
Assuming a very small prevalence in the endemic regime, i.e. $\rho_{I}^{bottleneck}\ll 1$, we can estimate $\rho_{I}^{bottleneck}$ as
$\rho_{I}^{bottleneck}\simeq\frac{2\mathcal{R}_{0}\widetilde{D}_{x}}{\mu\pi}\ .$ (11)
Appendix F: Influence of infectious period and population size on $\widetilde{D}^{C}_{x}$
Eq. (5) represents the critical evolution in the antigenic position $\widetilde{D}^{C}_{x}$ for the endemicity of viruses not evolving their infectiousness . While in the main manuscript we focus on showing the inverse dependence of $\widetilde{D}^{C}_{x}$ on $\mathcal{R}^{wt}_{0}$, here we are interested in exploring how the boundary of the endemic regime depends on the rest of parameters. To address this question, we first set the population size to $N=10000$, consequently fixing the constant $a=1/12$ (Fig. 4A), and represent in Fig. 8 the fraction of endemic realizations as a function of the reproduction number ${\cal R}_{0}^{wt}$ and the immune escape $\widetilde{D}_{x}$ for two additional values of the duration of the infectious window $\mu^{-1}$: $\mu^{-1}=8$ days (Fig. 8A) and $\mu^{-1}=10$ days (Fig. 8B). There we check that the theoretical estimation provided by Eq. 5 fairly captures the endemic regime for the three epidemic scenarios and confirm that the critical value $\widetilde{D}^{C}_{x}$ increases as the infectious period is reduced.
To estimate finite size effects on the virus endemicity, we represent in Figure 9 how the fraction of endemic realizations changes a function of the immune escape $\widetilde{D}_{x}$ and the population size $N$ for two pathogens differing in their reproduction number: $\mathcal{R}_{0}^{wt}=3$ (Fig. 9A) and $\mathcal{R}_{0}^{wt}=4$ (Fig. 9B). In both panels, we clearly show how the critical immune escape for endemicity decreases as populations become larger. Considering Eq. (5), this result reveals an inverse dependence of the scaling factor $a$ with the population size $N$. Namely, for small population sizes, a greater epidemic prevalence at the bottleneck is needed to prevent the extinction of epidemic outbreaks. Interestingly, the curves seem to collapse for large population sizes ($N\simeq 20000$), thus highlighting that $a(N)$ reaches an asymptomatic value for large population. The latter fact highlights that finite size effects are less relevant in large populations, for which transitions between endemic and non-endemic outbreaks are mainly driven by the interplay between epidemiological and evolutionary time scales.
Figure 8: Endemicity of viruses under deterministic evolution. A)-B): Fraction of endemic realizations $f_{endemic}$ as a function of the basic reproduction number of the wild-type variant $\mathcal{R}^{wt}_{0}$ and the speed of evolution in the antigenic space $\widetilde{D}_{x}$. The white solid line shows the theoretical estimation of the critical immunity escape value $\widetilde{D}^{C}_{x}$ delimiting the region $f_{endemic}=1$. Such quantity is obtained by setting $a=1/12$ and $b=5$ in Eq. 4 of the main text. The values considered for the duration of the infectious windows are: (A) $\mu^{-1}=10$ and (B) $\mu^{-1}=8$ days. In all panels, the speed of evolution in infectiousness is set to $\widetilde{D}_{\lambda}=0$. Moreover, the fraction of endemic realizations by performing 500 epidemic outbreaks and computing those persisting in the population after $t=1000$ days. The rest of model parameters are the same as in Fig. 2 of the main manuscript.
Figure 9: Finite size effects on the endemicity of viruses under deterministic evolution. Endemicity $f_{endemic}$ of the virus as a function of the basic reproduction number of the speed of evolution in the antigenic space $\widetilde{D}_{x}$. Two different values for the reproduction number of the wild-type variant are considered: (A) $\mathcal{R}_{0}^{wt}=3$ and (B) $\mathcal{R}_{0}^{wt}=4$. In all panels, line color denotes the population size NN and the speed of evolution in infectiousness is set to $\widetilde{D}_{\lambda}=0$ and the rest of epidemiological parameters are the same as in Fig. 2 of the main manuscript.
Appendix G: Time for an epidemic peak in a SIR model
@@
For a well-mixed population with constant size, the time to the epidemic peak $t_{peak}$ in the standard SIR model can be approximated as [39]:
@@
where $i_{0}$ refers to the fraction of population initially infected. This expression represents a first-order approximation to the exact computation of the epidemic peak in the standard SIR model. Nonetheless, it provides reasonably accurate estimates of the peak in most of the epidemic scenarios [39]. In our model, the latter expression is used to obtain a characteristic time scale during which viruses can evolve their transmissibility to avoid getting extinct. Consequently, note that the former expression comes with intrinsic limitations, as it assumes a fixed value of the infectiousness $\lambda$ and not a evolving quantity over time. Therefore, we should expect predictions to be misaligned with the actual epidemic peak for quickly evolving viruses with low initial infectiousness, i.e. $\mathcal{R}_{0}\simeq 1$}. In that region, the slow propagation of the virus leads to a substantial evolution in the infectiousness of the virus throughout the first epidemic wave, which in turn produces a much shorter epidemic wave than the one predicted by Eq. (12).
Figure 10: Non-monotonic behavior of epidemiological indicators at the bottleneck with ${\cal R}_{0}^{wt}$. A): Time evolution of number of infected individuals II for endemic epidemic outbreaks. B): Distribution of the minimum number of infected individuals at the epidemic bottleneck $I^{bottleneck}$ across endemic realizations as a function of the basic reproduction number of the wild-type variant ${\cal R}_{0}^{wt}$. C): Time evolution of effective infectiousness $\lambda_{eff}$ for endemic epidemic outbreaks. D): Distribution of the effective infectiousness at the epidemic bottleneck $\lambda^{bottleneck}_{eff}$ across endemic realizations as a function of ${\cal R}_{0}^{wt}$. In all panels, color encodes ${\cal R}_{0}^{wt}$. In panels (A) and (C), thick lines show average values whereas single lines correspond to single realizations. in panels (B) and (D), the grey dot denotes the average across endemic realizations whereas the whiskers of the boxplot denote IQR. In all the panels, the speeds of constant evolution have been set to $\widetilde{D}_{x}=0.003$ and $\widetilde{D}_{\lambda}=0.00004$ respectively. The rest of epidemiological parameters are the same as in Fig. 2
Appendix H: Non-monotonic behavior of epidemiological parameters in the bottleneck
In the main text, we find analytically an estimation of the critical speed of (constant) antigenic evolution $\widetilde{D}_{x}^{C}$ for a virus becoming endemic. This mathematical expression captures the non-monotonic behavior of viral endemicity with the infectiousness of the wild-type variant ${\cal R}_{0}^{wt}$. In this section, we explore whether there is a fingerprint of this non-monotonic behavior in the infection and evolution curves obtained from the simulations. Assuming $\widetilde{D}_{x}=0.003$ and $\widetilde{D}_{\lambda}=0.00004$, we present in Fig. 10A, the epidemic curves obtained for different values of ${\cal R}_{0}^{wt}$. One salient feature of these curves is that the prevalence at the epidemic bottleneck strongly varies with ${\cal R}_{0}^{wt}$. We characterize this phenomenon in Fig. 10B, where we show that there exists a non-monotonic behavior of the size of infected population with ${\cal R}_{0}^{wt}$. The latter makes outbreaks generated by viruses with intermediate infectiousness more likely to become extinct due to stochastic fluctuations.
In addition to the epidemic prevalence at the bottleneck, we also define here the effective infectiousness of the virus $\lambda_{eff}(t)$ as the product between the infectiousness of the circulating variant $\lambda_{c}$ and the fraction of population exposed of the disease at time $t$. Taking into account that all the infected population share the same strain under constant population, this indicator reads:
$\lambda_{eff}(t)=\sum_{i=1}^{N}\lambda_{c}(t)\left[\delta_{\Gamma_{i}(t),S}+\delta_{\Gamma_{i}(t),R}\left(1-e^{-(x_{c}(t)-x_{i})}\right)\right]\ ,$ (13)
where $\delta$ denotes the Kronecker delta and $\Gamma_{i}(t)$ the state of the agent $i$ at time $t$ during a given outbreak. Fig. 10C shows the evolution of $\lambda_{eff}$ for multiple outbreaks generated by viruses differing in ${\cal R}_{0}^{wt}$. We see that the model with constant evolution retrieves the universal long-term behavior of the evolution curves observed in the original model. Focusing on the epidemic bottleneck, Fig. 10D reveals that the effective infectiousness of the virus at the bottleneck also display a non-monotonic behavior with ${\cal R}_{0}^{wt}$. As stated in the manuscript, this arises from the trade-off between the original infectiousness of the viruses and the increments accumulated in their first epidemic waves. Taken together, our results show how viruses with ${\cal R}_{0}^{wt}$ are more likely to disappear due to the strong population bottlenecks after the first epidemic wave and an insufficient evolved effective infectiousness to create a new secondary outbreak.
Code and data availability statements
No new data have been generated in this article. Code needed to reproduce the results is publicly available at [68].